# Reconfigurable Intelligent Surface Aided Vehicular Edge Computing: Joint Phase-shift Optimization and Multi-User Power Allocation

### v1

@@ Abstract
Vehicular edge computing (VEC) is an emerging technology with significant potential in the field of internet of vehicles (IoV), enabling vehicles to perform intensive computational tasks locally or offload them to nearby edge devices. However, the quality of communication links may be severely deteriorated due to obstacles such as buildings, impeding the offloading process. To address this challenge, we introduce the use of Reconfigurable Intelligent Surfaces (RIS), which provide alternative communication pathways to assist vehicular communication. By dynamically adjusting the phase-shift of the RIS, the performance of VEC systems can be substantially improved. In this work, we consider a RIS-assisted VEC system, and design an optimal scheme for local execution power, offloading power, and RIS phase-shift, where random task arrivals and channel variations are taken into account. To address the scheme, we propose an innovative deep reinforcement learning (DRL) framework that combines the Deep Deterministic Policy Gradient (DDPG) algorithm for optimizing RIS phase-shift coefficients and the Multi-Agent Deep Deterministic Policy Gradient (MADDPG) algorithm for optimizing the power allocation of vehicle user (VU). Simulation results show that our proposed scheme outperforms the traditional centralized DDPG, Twin Delayed Deep Deterministic Policy Gradient (TD3) and some typical stochastic schemes.
@@ I Introduction
With the rapid development of internet of vehicles (IoV), vehicle edge computing (VEC), as a cutting-edge technology, is becoming an important solution to support real-time vehicle computing [1, 2]. Vehicle devices are usually equipped with limited processors and memory, which limits their capability to process complex algorithms and large data [3]. For example, high-precision real-time positioning services or driver assistance systems that are relied on real-time video analysis require significant computational resources, which may be beyond the capabilities of commonly deployed in-vehicle computing resources [4, 5]. However, VEC allows vehicles to offload tasks to edge servers with abundant computing and storage resources so that various functions of the vehicle can be achieved [6].
@@
Despite extensive research in VEC, challenges remain, particularly in urban environments where communication between vehicle users (VUs) and base stations (BSs) can be obstructed by buildings and other obstacles. These factors lead to attenuation and multipath effects on the wireless signals, which in turn affect the stability of the communication links and the transmission rate [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. In addition, communication between vehicles and infrastructure needs to be supported by sufficient bandwidth, especially for some applications requiring real-time updates, such as traffic flow monitoring or cooperative driving between vehicles. In high-density urban areas or during peak traffic hours, bandwidth may be limited, which can affect the timely transmission and processing of data for each vehicle [20, 21, 22, 23, 24, 25, 26].
@@
Recently, Reconfigurable Intelligent Surface (RIS) technology has demonstrated significant advantages in enhancing network quality by providing additional communication pathways [35]. RIS can dynamically adjust phase-shift to improve communication link quality, thereby enabling better task offloading to edge computing servers [36, 37]. Therefore, the integration of RIS into VEC systems is a research avenue that deserves to be explored extensively.
@@
However, optimizing the phase-shift of RIS is constrained by hardware conditions, which restrict the phase-shift to a finite set of values. In [38], He et al. have tackled the RIS phase-shift optimization problem using algorithms like Block Coordinate Descent (BCD), Alternating Optimization (AO), and Semi-definite Relaxation (SDR). However, these algorithms above have a high time complexity and make the RIS phase-shift coefficients adjustment be difficult in time [39]. In parallel, artificial intelligence (AI) has been applied to various engineering fields after rapid development. Deep Reinforcement Learning (DRL), as a type of AI, integrates deep learning (DL) and reinforcement learning (RL) methods, and becomes a promising technique that learns through the interaction of an intelligent body with its environment [40, 41, 42, 43]. It can effectively and quickly solve some non-convex problems with high complexity [27, 28, 29, 30, 31, 32, 33, 34].
@@
There are a number of challenges to leveraging RIS in vehicular communication environments. Firstly, the scenarios are complex and multiple VUs may exist, which can put pressure on training. Second, due to the high-speed mobility of vehicles, their locations are constantly changing, so the distances between the RIS and the VU, and between the VU and the BS, change over time, which affects the quality of the link between them [44]. Thirdly, the task arrival of each VU is random in practice. Fourthly, phase-shift optimization of RIS is complex, since RIS and VU are not the same type of agents, they may have different actions, That means we have to overcome the phase-shift optimization training of RIS first. Fifth, we should not only consider the power allocation problem of each VU, but also consider the buffer size at the same time. Sixth, how to combine the RIS phase-shift coefficient optimization with the VU power allocation problem is a challenging problem.
@@
The joint optimization of RIS phase-shift and power allocation for multiple VUs is too complex, and centralized algorithms can not work effectively in such dynamic settings [8]. In order to solve the problems mentioned above, we propose a novel power allocation and phase-shift optimization scheme for RIS-assisted vehicular communication. By using the modified Multi-Agent Deep Deterministic Policy Gradient (MADDPG) algorithm for VU power allocation and the Deep Deterministic Policy Gradient (DDPG) algorithm for RIS phase-shift optimization. The contributions of this work can be summarized as follows [^1] The source code has been released at: https://github.com/qiongwu86/DDPG-RIS-MADDPG-POWER:
@@
    A system model enhance VEC with RIS technique is proposed where the vehicle is in motion, the communication link is affected by obstacles, resulting in degradation of the link quality, and the VU ensures the capability of edge computing by offloading the tasks to the edge servers using RIS.
@@
    We investigate the joint RIS phase-shift coefficient optimization and VU power allocation in RIS-assisted the VEC. It is formulated as an optimization problem with the objective of minimizing the power consumption and buffer length of the VUs. However, due to various constraints, this is an intractable mixed integer non-convex problem.
  • 3)
    We decompose the optimization problem into two sub-problems, firstly, RIS is used as an auxiliary tool which aims to maximize the information transfer rate of the VUs by adjusting the phase-shift coefficients, so we first optimize the phase-shift coefficients of RIS by a centralized DDPG algorithm trained to obtain the phase-shift optimization model. Secondly, for the power allocation of VUs, we consider multiple users and use the modified MADDPG algorithm to solve the power allocation problem.
@@ II-A DRL-based Offloading Scheme for MEC
In [45], Liu et al. explored vehicle-assisted MEC networks and proposed a reinforcement learning-based offloading scheme to optimize task offloading and resource allocation. In [46], Huang et al. proposed a Deep Reinforcement Learning-based Online Offloading framework (DROO) for solving the task offloading problem in wireless-powered MEC networks. A expandable solution is achieved through Deep Q-Networks (DQN), which obviates the need to solve joint optimization problems and significantly reduces the computational complexity. In [47], Zhan et al. explored the task offloading problem in VEC networks and proposed an offloading scheduling method based on Deep Q Networks (DQN) to optimize task scheduling and resource allocation. In [48], Zheng et al. combined wireless power transfer (WPT) and multi-access edge computing (MAEC) to overcome energy and computational limitations in wireless devices. They proposed a DRL-based algorithm to minimize total computational delay (TCD) in WP-MEC networks by optimizing offloading decisions, WPT duration, and transmission duration, achieving near-optimal performance with low computational complexity in fast-fading channels. In [49], Alam et al. addressed computational offloading in high-mobility Internet of Vehicles (IoV) environments, ensuring latency, energy, and cost requirements. They proposed a cooperative three-tier decentralized Vehicle Assisted MAEC network and applied the Multi-Intelligent Deep Reinforcement Learning based Hungarian Algorithm (MADRLHA) to optimize dynamic task offloading, enhancing network performance and robustness. In [50], Liu et al. examined vehicular task offloading systems using reinforcement learning (RL) and DRL techniques, focusing on edge cloud (EC) and vehicular cloud (VC) computing. They categorized existing solutions involving MEC servers, nearby vehicles, and hybrid MEC (HMEC), summarizing lessons learned and outlining challenges for future research. In [51], Zhu et al. proposed a decentralized DRL framework to optimize power allocation in multi-input multi-output and non-orthogonal multiple access (MIMO-NOMA) VEC systems, addressing uncertainties in channel conditions and stochastic task arrivals. In [52], Tang et al. addressed the challenge of dynamic and efficient offloading decisions in VEC environments by proposing a Dynamic Framing Offloading algorithm based on Double Deep Q-Network (DFO-DDQN). It aims to minimize total delay and waiting time for compute-intensive tasks from multiple moving vehicles by optimizing offloading decisions at the subtask level in a frame-based model. In [53], Liu et al. explored collaborative task computing and dynamic resource allocation in VEC using an asynchronous deep reinforcement learning algorithm. It aims to optimize task offloading and resource management across vehicles, edge servers, and cloud environments. In [54], Ju et al. proposed NOMA-assisted secure offloading in VEC networks. It optimizes transmit power, resource allocation, and jammer vehicle selection to minimize energy consumption while meeting computation delay constraints. Using an A3C learning algorithm, the Energy-Efficiency Secure Offloading (EESO) scheme adapts dynamically to VEC environments, improving energy efficiency and ensuring security with low computation delay. In [55], Wu et al. presented a 5G-based vehicle-aware MAEC network, and proposed a deep reinforcement learning-based joint computation offloading and task migration optimization algorithm to optimize computation offloading and task migration. Experimental results demonstrated that it effectively reduces task delays, energy consumption, and improves resource utilization compared to alternative policies.
@@ II-B RIS-assisted Offloading Scheme for MEC
In [56], Zhang et al. introduced large-scale model (LM) techniques in IoV to tackle coverage challenges in dynamic wireless networks. They proposed a hybrid beamforming method using DRL to optimize beamforming with RIS and BS, demonstrating improved system throughput and coverage. In [57], Zhang et al. explored NOMA techniques with RIS in IoV to enhance signal transmission. They proposed power allocation schemes optimizing channel gain and data rate by adjusting RIS reflector units, enhancing system capacity and ensuring quality of service (QoS). In [58], Muhammad et al. investigated Zero-Energy RIS (ze-RIS) for vehicular computational offloading in 6G networks. They proposed a DRL-driven RIS-assisted energy-efficient task offloading (DREEO) scheme to save energy and time through power and offloading mechanism control. In [59], Chen et al. introduced MEC and RIS in IoT for low-latency task offloading. They developed an MEC-NOMA framework and RIS-assisted channel allocation algorithms to optimize offloading delays and transmission rates. In [60], Mao et al. applied RIS techniques in wirelessly powered MEC networks to improve energy efficiency and task offloading. They proposed an iterative algorithm to optimize upstream and downstream beamforming, transmission power, RIS time allocation, and IoT device computation frequency for maximizing total computation bits. In [61], Dai et al. integrated RIS with edge computing in 6G networks to enhance wireless communication efficiency and support low-latency applications. It introduces a deep reinforcement learning-based computation offloading scheme to minimize mobile device offloading latency, demonstrating improved data rates and reduced task execution times through RIS-aided enhancements. In [62], Hazarika et al. investigated an IoV network integrating MAEC servers at BSs with RISs for enhanced uplink and downlink transmissions. It employs a MA-DRL framework using Markov game theory to optimize task offloading decisions, improving network utility and communication quality significantly over existing methods. In [63], Xu et al. explored the integration of RIS into MEC networks, focusing on maximizing the sum computation rate. It proposes a DRL-based approach, specifically employing the TD3 algorithm, to optimize RIS phase shifts and energy partition strategies for Internet of Things devices (IoTDs). In [64], Ning et al. developed a DRL framework to jointly optimize computational edge servers, RIS deployment, and beamforming matrices in VEC networks, aiming to maximize the weighted sum throughput of Vehicular User Equipments (VUEs) while constraining latency within defined thresholds.
As mentioned above, in terms of designing RIS-assisted VEC power allocation and phase-shift optimization, this is the first time that we shall decompose this problem into two sub-problems and then find their solutions by DRL.
@@ III-A Scenario
As shown in Fig. 1, we consider a RIS-assisted VEC network with multiple users, where the BS is with $M$ antennas and a VEC server, and the VU has a single antenna. Since the VU has limited computational resources, it can offload some tasks to the VEC server for processing. When the VU’s link is blocked, the quality of its communication link with the BS is degraded, in which case the RIS with $N$ reflective elements can help the VU offload tasks to the edge device (It assumes that two links existing simultaneously, i.e., the VU communicates with the BS via the RIS and the VU communicates directly with the BS). Let us consider an urban road environment where the RIS is deployed on a building on one side of the road, where there are $K$ VUs, represented in the set as $\mathcal{K}=\left\{{1,2,\cdots,K}\right\}$. The VUs have the flexibility to offload tasks for processing or local execution, and in addition, we divide the task processing into $T$ equal discrete time slots $\Delta t$, denoted by $\mathcal{T}=\left\{{1,2,\cdots,T}\right\}$. At each time slot $t$ in $T$, the vehicle randomly generates a task whose arrival follows a Poisson distribution with an arrival rate of $\eta$. Note that we consider quasi-static scenarios, i.e., the channel conditions remain constant within a time slot but may change between neighboring time slots.

### v2

@@ Abstract
Vehicular edge computing (VEC) is an emerging technology with significant potential in the field of internet of vehicles (IoV), enabling vehicles to perform intensive computational tasks locally or offload them to nearby edge devices. However, the quality of communication links may be severely deteriorated due to obstacles such as buildings, impeding the offloading process. To address this challenge, we introduce the use of reconfigurable intelligent surface (RIS), which provide alternative communication pathways to assist vehicle communication. By dynamically adjusting the phase-shift of the RIS, the performance of VEC systems can be substantially improved. In this work, we consider a RIS-assisted VEC system, and design an optimal scheme for local execution power, offloading power, and RIS phase-shift, where random task arrivals and channel variations are taken into account. To address the scheme, we propose an innovative deep reinforcement learning (DRL) framework that combines the deep deterministic policy gradient (DDPG) algorithm for optimizing RIS phase-shift coefficients and the multi-agent DDPG (MADDPG) algorithm for optimizing the power allocation of vehicle user (VU). Simulation results show that our proposed scheme outperforms the traditional centralized DDPG, twin delayed deep deterministic policy gradient (TD3) and some typical stochastic schemes.
@@ I Introduction
With the rapid development of internet of vehicles (IoV), vehicular edge computing (VEC), as a cutting-edge technology, is becoming an important solution to support real-time vehicle computing [1, 2, 3]. Vehicle devices are usually equipped with limited processors and memory, which limits their capability to process complex algorithms and large data [4, 5]. For example, high-precision real-time positioning services or driver assistance systems that are relied on real-time video analysis require significant computational resources, which may be beyond the capabilities of commonly deployed in-vehicle computing resources [6, 7, 8]. However, VEC allows vehicles to offload tasks to edge servers with abundant computing and storage resources so that various functions of the vehicle can be achieved [9, 10, 11].
@@
Despite extensive research in VEC, challenges remain, particularu in urban environments where communication between vehicle users (VUs) and base stations (BSs) can be obstructed by buildings and other obstacles [12, 13, 14]. These factors lead to attenuation and multi-path effects on the wireless signals, which in turn affect the stability of the communication links and the transmission rate [15, 16]. In addition, communication between vehicles and infrastructure needs to be supported by sufficient bandwidth, especially for some applications requiring real-time updates, such as traffic flow monitoring or cooperative driving between vehicles [17, 18, 19]. In high-density urban areas or during peak traffic hours, bandwidth may be limited, which can affect the timely transmission and processing of data for each vehicle [20].
@@
Recently, reconfigurable intelligent surface (RIS) technology has demonstrated significant advantages in enhancing network quality by providing additional communication pathways [21, 22, 23]. RIS can dynamically adjust phase-shift to improve communication link quality, thereby enabling better task offloading to edge computing servers [24, 25]. Therefore, the integration of RIS into VEC systems is a research avenue that deserves to be explored extensively.
@@
However, optimizing the phase-shift of RIS is constrained by hardware conditions, which restrict the phase-shift to a finite set of values. In [26], He et al. have tackled the RIS phase-shift optimization problem using algorithms like Block Coordinate Descent (BCD), Alternating Optimization (AO), and Semi-definite Relaxation (SDR). However, these algorithms above have a high time complexity and make the RIS phase-shift coefficients adjustment be difficult in time [27]. In parallel, artificial intelligence (AI) has been applied to various engineering fields after rapid development [28, 29, 30, 31, 32, 33]. Deep Reinforcement Learning (DRL), as a type of AI, integrates deep learning (DL) and reinforcement learning (RL) methods, and becomes a promising technique that learns through the interaction of an intelligent body with its environment [34, 35, 36, 37, 38, 39]. It can effectively and quickly solve some non-convex problems with high complexity.
@@
There are a number of challenges to leveraging RIS in vehicular communication environments. Firstly, the scenarios are complex and multiple VUs may exist, which can put pressure on training. Second, due to the high-speed mobility of vehicles, their locations are constantly changing, so the distances between the RIS and the VU, and between the VU and the BS, change over time, which affects the quality of the link between them [40]. Thirdly, the task arrival of each VU is random in practice. Fourthly, phase-shift optimization of RIS is complex, since RIS and VU are not the same type of agents, they may have different actions, That means we have to overcome the phase-shift optimization training of RIS first. Fifth, we should not only consider the power allocation problem of each VU, but also consider the buffer length at the same time. Sixth, how to combine the RIS phase-shift coefficient optimization with the VU power allocation problem is a challenging problem.
@@
The joint optimization of RIS phase-shift and power allocation for multiple VUs is too complex, and centralized algorithms can not work effectively in such dynamic settings. In order to solve the problems mentioned above, we propose a novel power allocation and phase-shift optimization scheme for RIS-assisted VEC system. By using the modified multi-agent deep deterministic policy gradient (MADDPG) algorithm for VU power allocation and the deep deterministic policy gradient (DDPG) algorithm for RIS phase-shift optimization. The contributions of this work can be summarized as follows [^1] The source code has been released at: https://github.com/qiongwu86/DDPG-RIS-MADDPG-POWER:
@@
    A system model is proposed that enhances VEC communication using RIS technology. In this model, as the vehicle moves, communication links are impacted by obstacles, leading to reduced link quality. To address this, the VU leverages RIS to improve the quality of communication links and ensures efficient edge computing by offloading tasks to edge servers.
@@
    We investigate the joint optimization of RIS phase-shift coefficients and VU power allocation in RIS-assisted VEC systems. This problem is formulated with the objective of minimizing both power consumption and buffer length for the VUs. However, due to various constraints, it presents as a challenging mixed-integer non-convex problem.
    Since the offloading power and the local execution power are vehicle actions and the phase shift coefficients are RIS actions, the vehicle and the RIS are not the same type of agent. So, we decompose the optimization problem into two sub-problems, firstly, RIS is used as an auxiliary tool which aims to maximize the information transmission rate of the VUs by adjusting the phase-shift coefficients, so we first optimize the phase-shift coefficients of RIS by a centralized DDPG algorithm trained to obtain the phase-shift optimization model. Secondly, for the power allocation of VUs, we consider multiple users and use the modified MADDPG algorithm to solve the power allocation problem where the obtained phase-shift coefficients of RIS in previous phase is adopted as the given part.
@@ II-A DRL-based Offloading Scheme for MEC
In [41], Liu et al. explored vehicle-assisted MEC networks and proposed a DRL-based offloading scheme to optimize task offloading and resource allocation. In [42], Huang et al. proposed a DRL-based online offloading framework (DROO) for solving the task offloading problem in wireless-powered MEC networks. A expandable solution is achieved through deep Q-networks (DQN), which obviates the need to solve joint optimization problems and significantly reduces the computational complexity. In [43], Zhan et al. explored the task offloading problem in VEC networks and proposed an offloading scheduling method based DQN to optimize task scheduling and resource allocation. In [44], Zheng et al. combined wireless power transfer (WPT) and multi-access edge computing (MAEC) to overcome energy and computational limitations in wireless devices. They proposed a DRL-based algorithm to minimize total computational delay (TCD) in MEC networks by optimizing offloading decisions, WPT duration, and transmission duration, achieving near-optimal performance with low computational complexity in fast-fading channels. In [45], Alam et al. addressed computational offloading in high-mobility IoV environments, ensuring latency, energy, and cost requirements. They proposed a cooperative three-tier decentralized vehicle assisted MAEC network and applied the multi-agent DRL based Hungarian algorithm (MADRLHA) to optimize dynamic task offloading, enhancing network performance and robustness. In [46], Liu et al. examined vehicular task offloading systems using DRL techniques, focusing on edge cloud (EC) and vehicular cloud (VC) computing. They categorized existing solutions involving MEC servers, nearby vehicles, and hybrid MEC (HMEC), summarizing lessons learned and outlining challenges for future research. In [47], Zhu et al. proposed a decentralized DRL framework to optimize power allocation in multi-input multi-output and non-orthogonal multiple access (MIMO-NOMA) VEC systems, addressing uncertainties in channel conditions and stochastic task arrivals. In [48], Tang et al. addressed the challenge of dynamic and efficient offloading decisions in VEC environments by proposing a dynamic framing offloading algorithm based on double DQN. It aims to minimize total delay and waiting time for compute-intensive tasks from multiple moving vehicles by optimizing offloading decisions at the subtask level in a frame-based model. In [49], Liu et al. explored collaborative task computing and dynamic resource allocation in VEC using an asynchronous DRL algorithm. It aims to optimize task offloading and resource management across vehicles, edge servers, and cloud environments. In [50], Ju et al. proposed NOMA-assisted secure offloading in VEC networks. It optimizes transmit power, resource allocation, and jammer vehicle selection to minimize energy consumption while meeting computation delay constraints. Using an asynchronous advantage actor-critic (A3C) learning algorithm, the energy-efficiency secure offloading (EESO) scheme adapts dynamically to VEC environments, improving energy efficiency and ensuring security with low computation delay. In [51], Wu et al. presented a 5G-based vehicle-aware MAEC network, and proposed a DRL-based joint computation offloading and task migration optimization algorithm to optimize computation offloading and task migration.
@@ II-B RIS-assisted Offloading Scheme for MEC
In [52], Zhang et al. introduced large-scale model techniques in IoV to tackle coverage challenges in dynamic wireless networks. They proposed a hybrid beamforming method using DRL to optimize beamforming with RIS and BS, demonstrating improved system throughput and coverage. In [53], Zhang et al. explored NOMA techniques with RIS in IoV to enhance signal transmission. They proposed power allocation schemes optimizing channel gain and data rate by adjusting RIS reflector units, enhancing system capacity and ensuring quality of service (QoS). In [54], Muhammad et al. investigated zero-energy RIS (ze-RIS) for VEC in 6G networks. They proposed a DRL-driven RIS-assisted energy-efficient task offloading (DREEO) scheme to save energy and time through power and offloading mechanism control. In [55], Chen et al. introduced MEC and RIS in IoT for low-latency task offloading. They developed an MEC-NOMA framework and RIS-assisted channel allocation algorithms to optimize offloading delays and transmission rates. In [56], Mao et al. applied RIS techniques in wirelessly powered MEC networks to improve energy efficiency and task offloading. They proposed an iterative algorithm to optimize upstream and downstream beamforming, transmission power, RIS time allocation, and IoT device computation frequency for maximizing total computation bits. In [57], Dai et al. integrated RIS with edge computing in 6G networks to enhance wireless communication efficiency and support low-latency applications. It introduces a DRL-based computation offloading scheme to minimize mobile device offloading latency, demonstrating improved data rates and reduced task execution times through RIS-aided enhancements. In [58], Hazarika et al. investigated an IoV network integrating MAEC servers at BSs with RISs for enhanced uplink and downlink transmissions. It employs a MA-DRL framework using Markov game theory to optimize task offloading decisions, improving network utility and communication quality significantly over existing methods. In [59], Xu et al. explored the integration of RIS into MEC networks, focusing on maximizing the sum computation rate. It proposes a DRL-based approach, specifically employing the TD3 algorithm, to optimize RIS phase-shift and energy partition strategies for IoT devices. In [60], Ning et al. developed a DRL framework to jointly optimize computational edge servers, RIS deployment, and beamforming matrices in VEC networks, aiming to maximize the weighted sum throughput of VU equipment while constraining latency within defined thresholds. In [61], Xu et al. studied MEC networks with RIS support to maximize the total computation rate by optimizing the energy allocation strategy for RIS phase-shift and IoT devices, and then solve it by the TD3 algorithm. In [62], Bin et al. focused on optimizing energy consumption in MEC systems using RIS, especially in the context of mobile user equipment (UEs). The goal is to minimize the energy consumption by jointly optimizing the discrete phases of the RIS, the transmit power of the UEs, the computational resource allocation, and the task offloading policy. To cope with this complex problem with time-series decisions and uncertainties, a Soft Actor-Critic (SAC) algorithm based on deep reinforcement learning is proposed. In [63], a RIS-assisted MEC system is investigated, aiming to minimize the energy consumption of hybrid access point (HAP) by optimizing the offloading strategy of the user, the active beamforming of the hybrid access point and the passive beamforming of the IRS. An optimization-driven hierarchical deep deterministic policy gradient (OH-DDPG) framework is proposed to learn to adjust the RIS’s policy via external DDPG and improve the learning efficiency in multi-agent scenarios.
Reviewing the above related studies, we find that although many advances have been made in RIS-assisted VEC, comprehensive studies that simultaneously consider offloading power, local execution power, and phase-shift control are still relatively lacking. This is mainly due to the existing challenges faced by RIS-assisted VEC. In this paper, this is the first time that the integration of power allocation and phase-shift control in the RIS-assisted VEC are decomposed into two sub-problems and solved by the DRL method.
@@ III-A Scenario
As shown in Fig. 1, we consider a RIS-assisted VEC network with multiple VUs, where the BS is with $M$ antennas and a VEC server, and the VU has a single antenna. Since the VU has limited computational resources, it can offload some tasks to the VEC server for processing. When the VU’s link is blocked, the quality of its communication link with the BS is degraded, in which case the RIS with $N$ reflective elements can help the VU offload tasks to the edge device (It assumes that two links existing simultaneously, i.e., the VU communicates with the BS via the RIS and the VU communicates directly with the BS). Let us consider an urban road environment where the RIS is deployed on a building on one side of the road, where there are $K$ VUs, represented in the set as $\mathcal{K}=\left\{{1,2,\cdots,K}\right\}$. The VUs have the flexibility to offload tasks for processing or local execution, and in addition, we divide the task processing into $T$ equal discrete time slots $\Delta t$, denoted by $\mathcal{T}=\left\{{1,2,\cdots,T}\right\}$. At each time slot $t$ in $T$, the VU randomly generates a task whose arrival follows a Poisson distribution with an arrival rate of $\eta$. Note that we consider quasi-static scenarios, i.e., the channel conditions remain constant within a time slot but may change between neighboring time slots.